Total revenue of a particular production firm which engages in production for short run is Rs. 200 million, total cost is Rs.240 million and total fixed cost is Rs. 40 million. In this situitation should this firm continue its production or not. Explain.


1
Expert's answer
2022-02-13T11:51:41-0500

As total revenue equals total variable costs TVC = TC - TFC = 240 - 40 = 200 = TR, then the firm produces at the shutdown point and may continue producing while it can cover its variable costs.